3. Use the Gordon growth model to estimate Microsoft’s current stock price. Assume next year’s dividend payment is $12.00, the appropriate discount rate is 6 percent, and the company’s profits are expected to grow by 2 percent annually.
Answer #1

Next year's dividend payment (D1) = $12

Discount rate (rate of return, r) = 6% or 0.06

Growth rate (g) = 2% or 0.02

Using Gordon growth model,

Calculate the current stock price -

Current stock price = D1/(r - g) = $12/(0.06 - 0.02) = $12/0.04 = $300

Thus,

The Microsoft's current stock price is $300.